OpenStudy (anonymous):

Perform the indicated operation: 21x^5-56x^4+28x^8/7x^4

OpenStudy (anonymous):

is that meant to be \[\frac{21x^5-56x^4+28x^8}{7x^4}\ ?\]

OpenStudy (anonymous):

Yes, it meant to be that way.

OpenStudy (anonymous):

You may show all the work

OpenStudy (anonymous):

If so, then: \[\frac{21x^5-56x^4+28x^8}{7x^4}=\frac{7x^4(3x-8+4x^4)}{7x^4}=4x^4+3x-8\]